OpenStudy (anonymous):

Convert 288° to radians.

OpenStudy (anonymous):

Solve 1/6h - 2/3 = -1/4 for h h= 2 1/2 h= -11/72 h= -5 1/2 h= 5/72

OpenStudy (anonymous):

has to be fraction of pi

OpenStudy (anonymous):

To change degrees to radians, multiply degrees by pi/180. So 288 degrees = 288 x pi/180 = 8pi/5 simplified.

OpenStudy (jdoe0001):

\(\bf radians=\cfrac{degrees^o\pi}{180}\)

OpenStudy (campbell_st):

for radians you need \[180^0 = \pi\] so \[1^o = \frac{\pi}{180}\] so you have \[288 \times 1 = 288 \times \frac{\pi}{180}\]

OpenStudy (anonymous):

thanks Easyaspi314